OpenStudy (anonymous):

Susan is creating a solution. A formula for the solution calls for 7/10 mL of water. Susan wants to make 2/7 less of the solution than the formula creates. How much water should Susan use? Express your answer as a fraction in simplest form.

OpenStudy (anonymous):

making 2/7 less means she's making 5/7 of the original amount, so she needs 5/7 of 7/10 mL. \[\frac{ 5 }{ 7 }\times \frac{ 7 }{ 10 }\]